# Quantum automorphism groups of 2-graphs

### v2

@@ 4. Computations
  • The $2$-graph in the example (b) can be seen to be isomorphic to the cartesian product of the $1$-graph with two vertices and a bidirected edge (say $\mathcal{G}$) with itself in the sense of Proposition 1.8 of [NYJM]. In that example, the quantum automorphism group of the $2$-graph is ${\rm Aut}^{+}(\mathcal{G})\otimes{\rm Aut}^{+}(\mathcal{G})$. We conjecture that if $\mathcal{G}_{1},\mathcal{G}_{2}$ are two $1$-graphs, then the quantum automorphism group of the $2$-graph $\mathcal{G}_{1}\times\mathcal{G}_{2}$ at least contains ${\rm Aut}^{+}(\mathcal{G}_{1})\otimes{\rm Aut}^{+}(\mathcal{G}_{2})$ as a quantum subgroup.
  • We believe that the idea in this paper can be adapted to define the notion of the quantum automorphism group of $k$-graphs in general as a $k$-graph can be described by similar combinatorial data of $k$-colored graphs (see [Hazzlewood1]). But we believe that the computations will get harder as $k$ increases.
  • The quantum automorphism group of a $2$-graph should naturally act on the corresponding higher rank graph $C^{\ast}$-algebra as it has already been observed in the literature in the case of $1$-graphs (see [Joardar1],[Joardar2],[Weber] for example).
  • The quantum automorphism group of a $2$-graph has natural corepresentations on the morphism spaces of the $2$-graph. Therefore our consideration might also shed some light on the corepresentation theory of the quantum automorphism groups of $2$-graphs.
@@
This theorem must be well-known and proof being very natural, is only sketched here. Suppose $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent. Therefore both $(\mathcal{G}_{i},\mathcal{H}_{i},\theta_{i})$ for $i=1,2$ share the common vertex set $V$; a bijection $T:V\rightarrow V$; $\mathcal{G}_{i},\mathcal{H}_{i}$ have adjacency matrices $A_{i},B_{i}$ respectively for $i=1,2$ such that
@@
Let’s construct a functor $\Phi$ from $\Lambda_{1}$ to $\Lambda_{2}$. We recall that ${\rm obj}(\Lambda_{i})$ are identified with the common vertex set $V$. We assign to each $i\in V$ the element $T(i)\in V$. So that we associate each object of $\Lambda_{1}$ to an object in $\Lambda_{2}$. Now we want to associate each morphism in $\Lambda_{1}^{(m,n)}$ to a morphism in $\Lambda_{2}^{(m,n)}$ for all $(m,n)\in\mathbb{N}^{2}$. For $(m,n)=(1,0)$, a morphism $f$ from an object $i$ to an object $j$ in $\Lambda_{1}$ is an edge (say $f$) in the graph $\mathcal{G}_{1}$ with $s(f)=i$ and $r(f)=j$. Since $TA_{2}=A_{1}T$, there is an edge say $\Phi(f)$ with $s(\Phi(f))=T(i)$ and $r(\Phi(f))=T(j)$. Hence we get the following commutative diagram where $f\in\Lambda_{1}^{(1,0)}$ and $\Phi(f)\in\Lambda_{2}^{(1,0)}$:
{tikzcd}
@@
Similarly one gets a commutative diagram involving morphisms of degree $(0,1)$. Now let $F$ be a morphism of degree $(1,1)$ with $s(F)=i$ and $r(F)=j$. By the unique factorization property, there are unique morphisms $f_{1},f_{2}$ of degree $(1,0)$ and $g_{1},g_{2}$ of degree $(0,1)$ such that $F=f_{1}\circ g_{1}=g_{2}\circ f_{2}$. It is clear that two morphisms $f,g$ of degree $(1,0)$ or $(0,1)$ are composable iff $\Phi(f)$ and $\Phi(g)$ are composable. Then we define $\Phi(F)$ to be either of $\Phi(f_{1})\circ\Phi(g_{1})$ or $\Phi(g_{2})\circ\Phi(f_{2})$. Using the $\theta$-compatibility and the fact that $\theta_{1}(f_{1},g_{1})=(g_{2},f_{2})$ we get that $\theta_{2}(\Phi(f_{1}),\Phi(g_{1}))=(\Phi(g_{2}),\Phi(f_{2}))$. This proves the well-definedness of $\Phi(F)$. Now step by step one can define $\Phi(F)\in\Lambda_{2}^{(m,n)}$ for any $F\in\Lambda_{1}^{(m,n)}$ and for any $(m,n)\in\mathbb{N}^{2}$ and one easily verifies that $\Phi$ is a functor. Similarly one can define a functor $\Phi^{-1}$ from $\Lambda_{2}$ to $\Lambda_{1}$ using the map $T^{-1}$ on the vertex set $V$ (i.e. on the level of objects) and it is easily verified that $\Phi^{-1}$ is the inverse of the functor $\Phi$. Thus an isomorphism between the categories $\Lambda_{1}$ and $\Lambda_{2}$ is established.
@@
For the converse direction, we start with two isomorphic $2$-graphs $\Lambda_{1}$ and $\Lambda_{2}$ with defining triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ respectively. Using the isomorphism of $\Lambda_{1},\Lambda_{2}$, one can see that there is a bijection between ${\rm obj}(\Lambda_{1})$ and ${\rm obj}(\Lambda_{2})$ which are finite sets by our assumption. So one can without loss of generality assume that $\mathcal{G}_{i},\mathcal{H}_{i}$ share the same vertex set $V$. We denote the functor from $\Lambda_{1}$ to $\Lambda_{2}$ by $\Phi$. Then there is a morphism of degree $(1,0)$ from object $v_{i}$ to $v_{j}$ i.e. there is an edge of the graph $\mathcal{G}_{1}$ say $f$ from the vertex $v_{i}$ to the vertex $v_{j}$Rightarrow there is a unique morphism $\Phi(f)$ from $\Phi(v_{i})$ to $\Phi(v_{j})$ i.e. there is an edge from the vertex $\Phi(v_{i})$ to the vertex $\Phi(v_{j})$. Equivalently if the permutation matrix corresponding to the bijection $\Phi$ is denoted by $T$, $TA_{2}=A_{1}T$ where $A_{i}$ are the vertex matrices of $\mathcal{G}_{i}$ for $i=1,2$. Similarly considering degree $(0,1)$ morphisms one gets $TB_{2}=B_{1}T$. Let $F$ be a morphism of degree $(1,1)$ in $\Lambda_{1}$ with unique factorization $f_{1}\circ g_{1}=g_{1}^{\prime}\circ f_{1}^{\prime}$. Then, as usual, identifying morphisms of degree $(1,0)$ and $(0,1)$ in $\Lambda_{1}$ with edges of graphs $\mathcal{G}_{1},\mathcal{H}_{1}$ respectively, we get $\theta_{1}(f_{1},g_{1})=(g_{1}^{\prime},f_{1}^{\prime})$. But by the definition of functor, $\Phi(F)=\Phi(f_{1})\circ\Phi(g_{1})=\Phi(g_{1}^{\prime})\circ\Phi(f_{1}^{\prime})$. Therefore from the unique factorization of $\Phi(F)$ in $\Lambda_{2}$, we obtain that $\theta_{2}(\Phi(f_{1}),\Phi(g_{1}))=(\Phi(g_{1}^{\prime}),\Phi(f_{1}^{\prime}))$ so that we get the required $\theta$ compatibility to ensure that the triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent. $\Box$
@@
Acknowledgement The first author is supported by SERB MATRICS grant no. MTR/2022/000515 (Govt. of India). The second author is partially supported by JC Bose Fellowship and grant of Prof Debashish Goswami, ISI Kolkata, awarded to him by SERB, DST ( Govt of India ). Both the authors would like to thank Dr. Arnab Mandal for some useful discussions.

### v3

@@ 5. Examples and computations II: Free product and free wreath product
  • The $2$-graph in the example (b) of Section $4$ can be seen to be isomorphic to the cartesian product of the $1$-graph with two vertices and a bidirected edge (say $\mathcal{G}$) with itself in the sense of Proposition 1.8 of [NYJM]. In that example, the quantum automorphism group of the $2$-graph is ${\rm Aut}^{+}(\mathcal{G})\otimes{\rm Aut}^{+}(\mathcal{G})$. We conjecture that if $\mathcal{G}_{1},\mathcal{G}_{2}$ are two $1$-graphs, then the quantum automorphism group of the $2$-graph $\mathcal{G}_{1}\times\mathcal{G}_{2}$ at least contains ${\rm Aut}^{+}(\mathcal{G}_{1})\otimes{\rm Aut}^{+}(\mathcal{G}_{2})$ as a quantum subgroup.
  • One would like to find a ‘non-trivial’ example of a pair of $2$-graphs (or equivalently triples) that are not classically isomorphic, but quantum isomorphic. We are working in this direction and we hope to report it in a forthcoming article soon. To produce a “trivial” example, one can simply take a pair of quantum isomorphic graphs which are not classicaly isomorphic say ${\cal G}_{1},{\cal G}_{2}$. Then the triples $({\cal G}_{1},{\cal G}_{1},{\rm id})$ and $({\cal G}_{2},{\cal G}_{2},{\rm id})$ will be quantum equivalent but not classicaly equivalent.
  • The quantum automorphism group of a $2$-graph should naturally act on the corresponding higher rank graph $C^{\ast}$-algebra as it has already been observed in the literature in the case of $1$-graphs (see [Joardar1, Joardar2, Weber] for example).
  • We believe that the notion of quantum isomorphism between $2$-graphs given in the Definition 5.1 can be linked to a perfect quantum strategy for winning a suitable non-local “$2$-graph isomorphism game”.
@@
This theorem must be well-known and the proof being very natural, is only sketched here. Suppose $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent. Therefore we may essentially assume that both $(\mathcal{G}_{l},\mathcal{H}_{l},\theta_{l})$ for $l=1,2$ share the common vertex set $V$; a bijection $T:V\rightarrow V$; $\mathcal{G}_{l},\mathcal{H}_{l}$ have adjacency matrices $A_{l},B_{l}$ respectively for $l=1,2$ such that
@@
Let’s construct a functor $\Phi$ from $\Lambda_{1}$ to $\Lambda_{2}$. We recall that ${\rm obj}(\Lambda_{l})$ are identified with the common vertex set $V$. We assign to each $i\in V$ the element $T(i)\in V$. So that we associate each object of $\Lambda_{1}$ to an object in $\Lambda_{2}$. Now we want to associate each morphism in $\Lambda_{1}^{(m,n)}$ to a morphism in $\Lambda_{2}^{(m,n)}$ for all $(m,n)\in\mathbb{N}^{2}$. For $(m,n)=(1,0)$, a morphism $f$ from an object $i$ to an object $j$ in $\Lambda_{1}$ is an edge (say $f$) in the graph $\mathcal{G}_{1}$ with $s(f)=i$ and $t(f)=j$. Since $TA_{2}=A_{1}T$, there is an edge say $\Phi(f)$ with $s(\Phi(f))=T(i)$ and $t(\Phi(f))=T(j)$. Similarly one gets an edge $\Phi(g)$ with $s(\Phi(g))=T(i)$ and $t(\Phi(g))=T(j)$ for a morphism $g$ of degree $(0,1)$ with $s(g)=i$ and $t(g)=j$. Now let $F$ be a morphism of degree $(1,1)$ with $s(F)=i$ and $t(F)=j$. By the unique factorization property, there are unique morphisms $f_{1},f_{2}$ of degree $(1,0)$ and $g_{1},g_{2}$ of degree $(0,1)$ such that $F=f_{1}\circ g_{1}=g_{2}\circ f_{2}$. It is clear that two morphisms $f,g$ of degree $(1,0)$ or $(0,1)$ are composable iff $\Phi(f)$ and $\Phi(g)$ are composable. Then we define $\Phi(F)$ to be either of $\Phi(f_{1})\circ\Phi(g_{1})$ or $\Phi(g_{2})\circ\Phi(f_{2})$. Using the $\theta$-compatibility and the fact that $\theta_{1}(f_{1},g_{1})=(g_{2},f_{2})$ we get that $\theta_{2}(\Phi(f_{1}),\Phi(g_{1}))=(\Phi(g_{2}),\Phi(f_{2}))$. This proves the well-definedness of $\Phi(F)$. Now, for any $(m,n)\in\mathbb{N}^{2}$, step by step one can define $\Phi(F)\in\Lambda_{2}^{(m,n)}$ for $F\in\Lambda_{1}^{(m,n)}$. One can easily verify that $\Phi$ is a functor. Similarly, one can define a functor $\Phi^{-1}$ from $\Lambda_{2}$ to $\Lambda_{1}$ using the map $T^{-1}$ on the vertex set $V$ (i.e. on the level of objects) and it can be easily verified that $\Phi^{-1}$ is the inverse of the functor $\Phi$. Thus an isomorphism between the categories $\Lambda_{1}$ and $\Lambda_{2}$ is established.
@@
For the converse direction, we start with two isomorphic $2$-graphs $\Lambda_{1}$ and $\Lambda_{2}$ with defining triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ respectively. Using the isomorphism of $\Lambda_{1}$ and $\Lambda_{2}$, one can see that there is a bijection between ${\rm obj}(\Lambda_{1})$ and ${\rm obj}(\Lambda_{2})$ which are finite sets by our assumption. So, without loss of generality, one can assume that $\mathcal{G}_{l},\mathcal{H}_{l}$ share the same vertex set $V$. We denote the functor from $\Lambda_{1}$ to $\Lambda_{2}$ by $\Phi$. Then there is a morphism of degree $(1,0)$ from object $v_{i}$ to $v_{j}$ i.e. there is an edge of the graph $\mathcal{G}_{1}$ say $f$ from the vertex $v_{i}$ to the vertex $v_{j}$Rightarrow there is a unique morphism $\Phi(f)$ from $\Phi(v_{i})$ to $\Phi(v_{j})$ i.e. there is an edge from the vertex $\Phi(v_{i})$ to the vertex $\Phi(v_{j})$. Equivalently if the permutation matrix corresponding to the bijection $\Phi$ is denoted by $T$, $TA_{2}=A_{1}T$ where $A_{l}$ are the vertex matrices of $\mathcal{G}_{l}$ for $l=1,2$. Similarly considering degree $(0,1)$ morphisms one gets $TB_{2}=B_{1}T$. Let $F$ be a morphism of degree $(1,1)$ in $\Lambda_{1}$ with unique factorization $f_{1}\circ g_{1}=g_{1}^{\prime}\circ f_{1}^{\prime}$. Then, as usual, identifying morphisms of degree $(1,0)$ and $(0,1)$ in $\Lambda_{1}$ with edges of graphs $\mathcal{G}_{1},\mathcal{H}_{1}$ respectively, we get $\theta_{1}(f_{1},g_{1})=(g_{1}^{\prime},f_{1}^{\prime})$. But by the definition of functor, $\Phi(F)=\Phi(f_{1})\circ\Phi(g_{1})=\Phi(g_{1}^{\prime})\circ\Phi(f_{1}^{\prime})$. Therefore from the unique factorization of $\Phi(F)$ in $\Lambda_{2}$, we obtain that $\theta_{2}(\Phi(f_{1}),\Phi(g_{1}))=(\Phi(g_{1}^{\prime}),\Phi(f_{1}^{\prime}))$ so that we get the required $\theta$ compatibility to ensure that the triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ are equivalent. ∎
@@
Acknowledgement. The first author is supported by SERB MATRICS grant no. MTR/2022/000515 (Govt. of India). The second author is partially supported by JC Bose Fellowship and grant of Prof Debashish Goswami, ISI Kolkata, awarded to him by SERB, DST (Govt of India). The authors would like to thank Dr. Arnab Mandal for some useful discussions. The authors would also like to thank the anonymous referee for their useful comments.